# The dual nature of GHZ9: coexisting AGN and star formation activity in a remote X-ray source at z=10.145

### v3

@@ 4 Physical and morphological properties
The detection of several rest-frame emission lines (Table 1) in GHZ9 provides an opportunity to assess its metallicity (Z), electron temperature (Te), nebular reddening (E(B-V)), nitrogen-to-oxygen (N/O), and carbon-to-oxygen (C/O) abundance ratios. Moreover from the photometry we assess its morphology and stellar masses.
Metallicity estimates are derived from the Ne3O2 ratio-based calibrations (Shi et al., 2007; Maiolino et al., 2008; Jones et al., 2015; Bian et al., 2018; Mingozzi et al., 2022; Curti et al., 2023), yielding Z [0.01, 0.1] Z⊙. This result aligns with recent studies of high redshift galaxies (e.g., Goulding et al., 2023; Hsiao et al., 2024; Carniani et al., 2025; Castellano et al., 2024; Maiolino et al., 2024b; Schouws et al., 2024), which show metal enrichment already at z 10. The ionization parameter is constrained using the C IV/C III] and the C IV EW relations (Mingozzi et al., 2022) and the Ne3O2 relation (Witstok et al., 2021), yielding log = [-1.90, -1.65].
We perform an AGN+SFG SED fit using the dale2014 module (Dale et al., 2014) from CIGALE (Boquien et al., 2019) and considering the available HST and JWST photometry (Fig. 2). We adopt a flexible star formation history (SFH) with two parametric components to take in consideration the presence of both old and young stellar populations. Namely, we include a “delayed” component of age 200 Myr and allow for a recent exponential burst with timescale 150 Myr and age between 5 and 100 Myr. We assume a Chabrier (2003) initial-mass function (IMF), a Calzetti et al. (2000) extinction law and we restrict metallicity and ionization parameter to the ranges estimated from the NIRSpec spectrum. Namely, the gas metallicity can be 2, 5 and 10% the solar value, while the stellar metallicity is fixed to 2% solar, and log can vary in the range [-1.90, -1.60] with 0.1 steps. The AGN component is parameterized by the AGN fraction (fAGN), defined as the ratio of the AGN luminosity to the total AGN and dust luminosities. We initially performed the fit with the AGN fraction set as a free parameter and obtained a best fit likelihood-weighted value of fAGN=0.30 (68% confidence region 0.01–0.6), resulting in a lens-corrected stellar mass of 4.9 108 M⊙. To better understand the impact of the AGN on the stellar mass estimate, we repeated the fit fixing fAGN at discrete values within the 0–0.6 range. The lens-corrected stellar mass was found to vary between 3.3 and 7.2 108 M⊙, for the maximum and null AGN contribution, respectively. The best-fit nebular reddening is E(B–V) = 0.32 0.18.
@@
We measured the electron temperature from the flux ratio between O III] 1661,66 and [O III] 4363 lines, correcting for dust attenuation using the Calzetti et al. (2000) law and the E(B-V) value derived from the SED fitting. Assuming an electron density of 103 cm-3, we used PyNeb to compute a conservative 2 (1) lower limit based on the corresponding lower bounds of the flux ratio. This yields Te = 2 104 K (4 104 K), in good agreement with recent results for high-redshift AGN (e.g., Tripodi et al., 2024, see their Fig. 4). The large uncertainty in E(B–V), combined with the limited spectral resolution that prevents decomposition of the O III] 1661,66 components, limits a more precise determination of Te.
@@
We use the measured line ratios N IV] /O III] and N III] /O III] and consider a range of electron densities ([103, 5 103, 104, 5 104, 105, and 5 105] cm-3) and temperatures ([1.5, 2, 2.5, 3] 104 K). We performed a Monte Carlo analysis by perturbing the observed fluxes by their corresponding uncertainties 1000 times. The resulting N/O values range from -0.08 (for ne= 5 105cm-3 and Te=15000 K) to 0.12 (for ne=103cm-3 and Te=30000 K), which are 6–9.5 times higher than the solar N/O value (log(N/O)⊙ = -0.86; Asplund et al., 2009).
@@
For the measured ranges of and log , the resulting ICF ranges from 1.10 to 1.32. Using these parameters, we derive log(C/O) values ranging from -0.96 (for ne= 5 105cm-3 and Te=15000 K) to -0.45 (for ne=103cm-3 and Te=30000 K), which are 0.2–0.65 times the solar C/O value (log(C/O)⊙ = -0.26; Asplund et al., 2009).
We note that the exceptionally high C III] EW = (48 5) Å in GHZ9 is matched by only one other source in the literature, UNCOVER-45924, a BLAGN at z=4.5 (Greene et al., 2024; Treiber et al., 2025). UNCOVER-45924 is associated with a secure [Ne V] detection, whereas GHZ9 lacks high-ionization lines ( 60 eV) such as [Ne IV] and [Ne V], for which we provide a 3 upper limit in Table 1. However, we find significant (S/N 3) detections of the C II and Si IV multiplets. As discussed in Maiolino et al. (2024b) for GNz11, these lines are commonly observed in AGN spectra (e.g. Wu & Shen, 2022). Higher resolution observations are needed to better constrain the multiplets and determine if we can detect a broad component from permitted lines.
@@ 5 Evidence of AGN emission from UV and optical line diagnostics
The detection of several rest-frame emission lines in GHZ9 (Table 1) provides an opportunity to assess whether its primary ionizing source is an AGN or stellar populations. To achieve this, we use a combination of UV emission line ratios and EW diagnostics from the literature (Feltre et al., 2016; Gutkin et al., 2016; Nakajima et al., 2018; Hirschmann et al., 2019). Specifically, we analyze the EW of key UV metal lines (C IV 1548,51 , O III] 1661,66 , N III] 1747,49 , C III] 1908) and their ratios to the He II 1640 recombination line, as these are sensitive indicators to the hardness of the ionizing radiation.
@@
We compare these measurements to the photoionization models by Nakajima & Maiolino (2022) (hereafter NM22), constructed with the CLOUDY code (Ferland et al., 2013). These models consider a variety of ionization sources, including SFG, AGN, Population III stars, and Direct Collapse black holes (DCBH) and span a broad range of physical parameters, including gas metallicities, ionization parameters, and gas densities (see also Nakajima et al., 2018). They also incorporate BPASS stellar population synthesis models (Eldridge et al., 2017). In our analysis, we restrict the comparison to the NM22 AGN and SFG grids that are consistent with the physical properties inferred for GHZ9. In particular, we explore models with -2 log -1.5, 10-4 Zgas 2 10-3, and nH=103cm-3. The models account for varying C/O and N/O abundance ratios as a function of metallicity (Dopita et al., 2006; Nakajima et al., 2018).
@@
We apply the same parameter constraints in log and Zgas when comparing GHZ9 to the photoionization models by Feltre et al. (2016) and Gutkin et al. (2016) (hereafter FG16). In this case, the SFG models are further restricted to 0.14 (C/O)/(C/O)⊙ 0.72. The nitrogen abundance follows the prescription of Groves et al. (2004) in their grid. To aid interpretation, we include the UV diagnostic demarcation lines from Hirschmann et al. (2019) (see also, Hirschmann et al., 2023), which distinguish between AGN-dominated, SFG-dominated, and composite sources. These diagnostics were designed to maximize AGN classification purity (approximately 90%) up to z=8, which corresponds to 180 Myr of galaxy evolution from GHZ9. Additionally, we employ the optical diagnostic diagram introduced by Mazzolari et al. (2024b), based on [O III] 4363 / H vs. [Ne III] 3869 / [O II] 3727,29. While [Ne III]/[O II] (Ne3O2) traces the ionization state of the interstellar medium, [O III] 4363 provides information about the electron temperature, thereby offering insight into the energy output of the ionizing source. We note that both the Hirschmann et al. (2019) and Mazzolari et al. (2024b) diagnostics were derived from a broader space than the one strictly matching GHZ9’s properties (see Table 1 in Hirschmann et al., 2017). Fig. 3 shows the position of GHZ9 in these diagnostic diagrams and compares it with other spectroscopically confirmed AGN candidates at z 8.5, including UNCOVER-20466 (z=8.50; Kokorev et al., 2023), CEERS-1019 (z=8.68; Larson et al., 2023a), GS-z9-0 (z=9.43; Curti et al., 2025), GNz11 (z=10.6; Bunker et al., 2023; Maiolino et al., 2024b), and GHZ2 (z=12.34; Castellano et al., 2024). For comparison we show the AGN and SFG photoionization models from NM22 and FG16.
@@
As discussed by Castellano et al. (2024) and Maiolino et al. (2024b), the interpretation of whether the observed high redshift AGN candidates, including GHZ9, are primarily AGN- or SF-dominated is strongly model-dependent. Here, we detail our analysis based on the diagnostic diagrams presented in Fig. 3. In the C III]/He II versus C IV/C III] and the C III]/He II versus O III]/He II diagnostics, GHZ9 lies in the composite region, consistent within 1 with both the explored AGN and SFG models predictions. Therefore, these diagnostics do not allow us to determine the dominant ionizing source powering the observed emission lines.
@@
In the optical [O III]/H versus [Ne III]/[O II] diagram, GHZ9 lies in the AGN-dominated region, though it is consistent within 2 with the composite region and with both AGN and SFG predictions by FG16. When using the NM22 models, however, only the AGN scenario is compatible with the observed emission lines and their uncertainties.
@@
Finally, using the nitrogen based diagnostics (C III]/He II versus N III]/He II and EW NIII] versus N III]/He II) the position of GHZ9 is compatible with SFG-dominated and composite regions. However the observed EW of N III] exceeds theoretical predictions by more than an order of magnitude. No AGN or SFG model explored in our analysis reproduces such extreme EW values. This suggests that these diagrams primarily reflect the N-enriched nature of GHZ9 rather than constraining its ionizing source. To our knowledge, no existing models are designed to distinguish AGN from star formation emission in N-enhanced sources.
![[Uncaptioned image]](x2.png)
![[Uncaptioned image]](x3.png)
![[Uncaptioned image]](x4.png)
![[Uncaptioned image]](x5.png)
![[Uncaptioned image]](x6.png)
![[Uncaptioned image]](x7.png)
@@ 6 Nature of the X-ray emission
We first assessed the reliability of the association between the X-ray detection and GHZ9. As discussed in Sect. 3, the Chandra X-ray counts have been measured within a circular region of 1′′ radius where two other NIRCam-detected objects are found. The sources are ID=29686 and ID=29852 from the catalog by Merlin et al. (2024), located 0.5′′ from GHZ9. The spectrum of ID=29686 was serendipitously observed in the same slit as GHZ9. Based on the H emission, the unresolved [OIII]4959, 5007, and [OII]3727, 3729 doublets, we determine the spectroscopic redshift to be zspec = 1.117 0.006, with a corresponding magnification = 1.23 from Bergamini et al. (2023). We further characterized this source to investigate, using the log ([OIII] / H) vs log (M/M⊙) diagnostic (MEx diagram, Juneau et al., 2011), whether it could be compatible with an AGN nature. From SED fitting analysis, we find its stellar mass to be log (M/M⊙) = 7.65 and a null E(B-V) measurement. The [OIII] and H fluxes were obtained assuming no dust correction and case B recombination with a density = 100 cm-3 and temperature = 10,000 K. The results indicate a star-forming nature, with the AGN scenario disfavored at a significance level of 4.5 .
Instead, ID=29852 is an ultrafaint source (29.7), whose best-fit model corresponds to a low-mass (log (M/M⊙) = 6.3) passive (sSFR10-11 yr-1) galaxy at zphot = 0.575. Assuming the X-ray emission to be associated with this object, the resulting BH mass would be 2.6 104 M⊙. In this scenario the galaxy would be hosting an intermediate-mass black hole (102–105 M⊙), a rare class of objects for which no direct identification has been obtained beyond the local Universe (e.g., Greene et al., 2020; Boorman et al., 2024).
We finally assessed whether stellar processes in ID=29686 and ID=29852 could account for the observed X-ray emission by using the SFR values derived from the SED fitting and applying the X-ray luminosity relation for star-forming galaxies from Eq. 14 in Lehmer et al. (2016). We find that the expected X-ray luminosities (1039 and 1037erg s-1, respectively) are more than three orders of magnitude lower than the observed ones (1042 and 1041erg s-1), effectively ruling out a stellar origin for the X-ray emission in both cases. Similarly, for GHZ9, we exclude the possibility that the X-ray emission is due to stellar processes, as the measured X-ray luminosity (Sect. 3) for an object with its stellar mass would require a SFR more than two orders of magnitude higher than what is estimated for this source.
@@
6.1 The super-massive black hole in GHZ9
The X-ray detection provides evidence that GHZ9 hosts an accreting SMBH. We derive a bolometric luminosity of 2.0 (0.6) erg s-1for = 2.3 ( = 1.8), assuming standard bolometric corrections for quasars (Duras et al., 2020). This bolometric luminosity is consistent with an alternative estimate derived from the continuum at 4400 Å (Duras et al., 2020), after correcting for the measured extinction value (see Sect. 4), thus pointing to a significant AGN contribution to the optical emission.
The bolometric luminosity derived implies that, if the BH in GHZ9 is radiating at its Eddington limit (fEdd = 1), its mass is 1.600.31 (0.480.09) , consistent with the value reported by Kovács et al. (2024). We note that the reported uncertainty on the BH mass only includes the 0.5–7 keV error counts. Additional uncertainties, due to the intrinsic scatter in the bolometric corrections, could be as high as a factor of 2 (see Figure 2 in Duras et al., 2020).

### v4

@@ 4 Physical and morphological properties
The detection of several rest-frame emission lines (Table 1) in GHZ9 provides an opportunity to assess its metallicity (Z), electron temperature (Te), nebular reddening (E(B-V)), nitrogen-to-oxygen (N/O), and carbon-to-oxygen (C/O) abundance ratios. Moreover, from the photometry we assess its morphology and stellar masses.
Metallicity estimates are derived from the Ne3O2 ratio-based calibrations (Shi et al., 2007; Maiolino et al., 2008; Jones et al., 2015; Bian et al., 2018; Mingozzi et al., 2022; Curti et al., 2023), yielding Z [0.01, 0.1] Z⊙. This result aligns with recent studies of high-redshift galaxies (e.g., Goulding et al., 2023; Hsiao et al., 2024; Carniani et al., 2025; Castellano et al., 2024; Maiolino et al., 2024b; Schouws et al., 2025), which show metal enrichment already at z 10. The ionization parameter is constrained using the C IV/C III] and the C IV EW relations (Mingozzi et al., 2022) and the Ne3O2 relation (Witstok et al., 2021), yielding log = [-1.90, -1.65].
We perform an AGN+SFG SED fit using the dale2014 module (Dale et al., 2014) from CIGALE (Boquien et al., 2019) and considering the available HST and JWST photometry (Figure 2). We adopt a flexible SFH with two parametric components to take in consideration the presence of both old and young stellar populations. Namely, we include a “delayed” component of age 200 Myr and allow for a recent exponential burst with timescale 150 Myr and age between 5 and 100 Myr. We assume a Chabrier (2003) initial-mass function (IMF), a Calzetti et al. (2000) extinction law and we restrict metallicity and ionization parameter to the ranges estimated from the NIRSpec spectrum. Namely, the gas metallicity can be 2%, 5%, and 10% the solar value, while the stellar metallicity is fixed to 2% solar, and log can vary in the range [-1.90, -1.60] with 0.1 steps. The AGN component is parameterized by the AGN fraction (fAGN), defined as the ratio of the AGN luminosity to the total AGN and dust luminosities. We initially performed the fit with the AGN fraction set as a free parameter and obtained a best-fit likelihood-weighted value of fAGN=0.30 (68% confidence region 0.01–0.6), resulting in a lens-corrected stellar mass of 4.9 108 M⊙. To better understand the impact of the AGN on the stellar-mass estimate, we repeated the fit fixing fAGN at discrete values within the 0–0.6 range. The lens-corrected stellar mass was found to vary between 3.3 and 7.2 108 M⊙, for the maximum and null AGN contribution, respectively. The best-fit nebular reddening is E(B–V) = 0.32 0.18.
@@
We measured the electron temperature from the flux ratio between O III] 1661,66 and [O III] 4363 lines, correcting for dust attenuation using the Calzetti et al. (2000) law and the E(B-V) value derived from the SED fitting. Assuming an electron density of 103 cm-3, we used PyNeb to compute a conservative 2 (1) lower limit based on the corresponding lower bounds of the flux ratio. This yields Te = 2 104 K (4 104 K), in good agreement with recent results for high-redshift AGN (e.g., Tripodi et al., 2024, see their Figure 4). The large uncertainty in E(B–V), combined with the limited spectral resolution that prevents decomposition of the O III] 1661,66 components, limits a more precise determination of Te.
@@
We use the measured line ratios N IV] /O III] and N III] /O III] and consider a range of electron densities ([103, 5 103, 104, 5 104, 105, and 5 105] cm-3) and temperatures ([1.5, 2, 2.5, 3] 104 K). We performed a Monte Carlo analysis by perturbing the observed fluxes by their corresponding uncertainties 1000 times. The resulting N/O values range from -0.08 (for ne= 5 105cm-3 and Te=15,000 K) to 0.12 (for ne=103cm-3 and Te=30,000 K), which are 6–9.5 times higher than the solar N/O value (log(N/O)⊙ = -0.86; Asplund et al., 2009).
@@
For the measured ranges of and log , the resulting ICF ranges from 1.10 to 1.32. Using these parameters, we derive log(C/O) values ranging from -0.96 (for ne= 5 105cm-3 and Te=15,000 K) to -0.45 (for ne=103cm-3 and Te=30,000 K), which are 0.2–0.65 times the solar C/O value (log(C/O)⊙ = -0.26; Asplund et al., 2009).
We note that the exceptionally high C III] EW = (48 5) Å in GHZ9 is matched by only one other source in the literature, UNCOVER-45924, a BLAGN at z=4.5 (Greene et al., 2024; Treiber et al., 2025). UNCOVER-45924 is associated with a secure [Ne V] detection, whereas GHZ9 lacks high-ionization lines ( 60 eV) such as [Ne IV] and [Ne V], for which we provide a 3 upper limit in Table 1. However, we find significant (S/N 3) detections of the C II and Si IV multiplets. As discussed in Maiolino et al. (2024b) for GN-z11, these lines are commonly observed in AGN spectra (e.g. Wu & Shen, 2022). Higher resolution observations are needed to better constrain the multiplets and determine if we can detect a broad component from permitted lines.
@@ 5 Evidence of AGN emission from UV and optical line diagnostics
The detection of several rest-frame emission lines in GHZ9 (Table 1) provides an opportunity to assess whether its primary ionizing source is an AGN or stellar populations. To achieve this, we use a combination of UV emission line ratios and EW diagnostics from the literature (Feltre et al., 2016; Gutkin et al., 2016; Nakajima et al., 2018; Hirschmann et al., 2019). Specifically, we analyze the EW of key UV metal lines (C IV 1548,51 , O III] 1661,66 , N III] 1747,49 , C III] 1908) and their ratios to the He II 1640 recombination line because they are sensitive indicators to the hardness of the ionizing radiation.
@@
We compare these measurements to the photoionization models by Nakajima & Maiolino (2022) (hereafter NM22), constructed with the CLOUDY code (Ferland et al., 2013). These models consider a variety of ionization sources, including SFG, AGN, Population III stars, and direct collapse black holes and span a broad range of physical parameters, including gas metallicities, ionization parameters, and gas densities (see also Nakajima et al., 2018). They also incorporate BPASS stellar population synthesis models (Eldridge et al., 2017). In our analysis, we restrict the comparison to the NM22 AGN and SFG grids that are consistent with the physical properties inferred for GHZ9. In particular, we explore models with -2 log -1.5, 10-4 Zgas 2 10-3, and nH=103cm-3. The models account for varying C/O and N/O abundance ratios as a function of metallicity (Dopita et al., 2006; Nakajima et al., 2018).
@@
We apply the same parameter constraints in log and Zgas when comparing GHZ9 to the photoionization models by Feltre et al. (2016) (hereafter F16) and Gutkin et al. (2016) (hereafter G16). In this case, the SFG models are further restricted to 0.14 (C/O)/(C/O)⊙ 0.72. The nitrogen abundance follows the prescription of Groves et al. (2004) in their grid. To aid interpretation, we include the UV diagnostic demarcation lines from Hirschmann et al. (2019) (see also, Hirschmann et al., 2023), which distinguish between AGN-dominated, SFG-dominated, and composite sources. These diagnostics were designed to maximize AGN classification purity (approximately 90%) up to z=8, which corresponds to 180 Myr of galaxy evolution from GHZ9. Additionally, we employ the optical diagnostic diagram introduced by Mazzolari et al. (2024b), based on [O III] 4363 / H versus [Ne III] 3869 / [O II] 3727,29. While [Ne III]/[O II] (Ne3O2) traces the ionization state of the interstellar medium, [O III] 4363 provides information about the electron temperature, thereby offering insight into the energy output of the ionizing source. We note that both the Hirschmann et al. (2019) and Mazzolari et al. (2024b) diagnostics were derived from a broader space than the one strictly matching GHZ9’s properties (see Table 1 in Hirschmann et al., 2017). Figure 3 shows the position of GHZ9 in these diagnostic diagrams and compares it with other spectroscopically confirmed AGN candidates at z 8.5, including UNCOVER-20466 (z=8.50; Kokorev et al., 2023), CEERS-1019 (z=8.68; Larson et al., 2023a), GS-z9-0 (z=9.43; Curti et al., 2025), GN-z11 (z=10.6; Bunker et al., 2023; Maiolino et al., 2024b), and GHZ2 (z=12.34; Castellano et al., 2024). For comparison we show the AGN and SFG photoionization models from NM22, F16, and G16.
@@
As discussed by Castellano et al. (2024) and Maiolino et al. (2024b), the interpretation of whether the observed high-redshift AGN candidates, including GHZ9, are primarily AGN- or SFG-dominated is strongly model-dependent. Here, we detail our analysis based on the diagnostic diagrams presented in Figure 3. In the C III]/He II versus C IV/C III] and the C III]/He II versus O III]/He II diagnostics, GHZ9 lies in the composite region, consistent within 1 with both the explored AGN and SFG models predictions. Therefore, these diagnostics do not allow us to determine the dominant ionizing source powering the observed emission lines.
@@
In the optical [O III]/H versus [Ne III]/[O II] diagram, GHZ9 lies in the AGN-dominated region, though it is consistent within 2 with the composite region and with both AGN and SFG predictions by F16 and G16. When using the NM22 models, however, only the AGN scenario is compatible with the observed emission lines and their uncertainties.
@@
Finally, using the nitrogen-based diagnostics (C III]/He II versus N III]/He II and EW NIII] versus N III]/He II) the position of GHZ9 is compatible with SFG-dominated and composite regions. However the observed EW of N III] exceeds theoretical predictions by more than an order of magnitude. No AGN or SFG model explored in our analysis reproduces such extreme EW values. This suggests that these diagrams primarily reflect the N-enriched nature of GHZ9 rather than constraining its ionizing source. To our knowledge, no existing models are designed to distinguish AGN from star formation emission in N-enhanced sources.
@@ 6 Nature of the X-ray emission
We first assessed the reliability of the association between the X-ray detection and GHZ9. As discussed in Section 3, the Chandra X-ray counts have been measured within a circular region of 1′′ radius where two other NIRCam-detected objects are found. The sources are ID=29686 and ID=29852 from the catalog by Merlin et al. (2024), located 0.5′′ from GHZ9. The spectrum of ID=29686 was serendipitously observed in the same slit as GHZ9. Based on the H emission, the unresolved [OIII]4959, 5007, and [OII]3727, 3729 doublets, we determine the spectroscopic redshift to be zspec = 1.117 0.006, with a corresponding magnification = 1.23 from Bergamini et al. (2023). We further characterized this source to investigate, using the log ([OIII] / H) versus log (M/M⊙) diagnostic (MEx diagram, Juneau et al., 2011), whether it could be compatible with an AGN nature. From SED fitting analysis, we find its stellar mass to be log (M/M⊙) = 7.65 and a null E(B-V) measurement. The [OIII] and H fluxes were obtained assuming no dust correction and case B recombination with a density = 100 cm-3 and temperature = 10,000 K. The results indicate a star-forming nature, with the AGN scenario disfavored at a significance level of 4.5 .
Instead, ID=29852 is an ultrafaint source (29.7), whose best-fit model corresponds to a low-mass (log (M/M⊙) = 6.3) passive (sSFR10-11 yr-1) galaxy at zphot = 0.575. Assuming the X-ray emission to be associated with this object, the resulting BH mass would be 2.6 104 M⊙. In this scenario the galaxy would be hosting an intermediate-mass BH (102–105 M⊙), a rare class of objects for which no direct identification has been obtained beyond the local Universe (e.g., Greene et al., 2020; Boorman et al., 2024).
We finally assessed whether stellar processes in ID=29686 and ID=29852 could account for the observed X-ray emission by using the SFR values derived from the SED fitting and applying the X-ray luminosity relation for SFGs from Equation (14) in Lehmer et al. (2016). We find that the expected X-ray luminosities (1039 and 1037erg s-1, respectively) are more than three orders of magnitude lower than the observed ones (1042 and 1041erg s-1), effectively ruling out a stellar origin for the X-ray emission in both cases. Similarly, for GHZ9, we exclude the possibility that the X-ray emission is due to stellar processes because the measured X-ray luminosity (Section 3) for an object with its stellar mass would require a SFR more than two orders of magnitude higher than what is estimated for this source.
@@
6.1 The Supermassive Black Hole in GHZ9
The X-ray detection provides evidence that GHZ9 hosts an accreting SMBH. We derive a bolometric luminosity of 2.0 (0.6) erg s-1for = 2.3 ( = 1.8), assuming standard bolometric corrections for quasars (Duras et al., 2020). This bolometric luminosity is consistent with an alternative estimate derived from the continuum at 4400 Å (Duras et al., 2020), after correcting for the measured extinction value (see Section 4), thus pointing to a significant AGN contribution to the optical emission.
The bolometric luminosity derived implies that, if the BH in GHZ9 is radiating at its Eddington limit (fEdd = 1), its mass is 1.600.31 (0.480.09) , consistent with the value reported by Kovács et al. (2024). We note that the reported uncertainty in the BH mass only includes the 0.5–7 keV error counts. Additional uncertainties, due to the intrinsic scatter in the bolometric corrections, could be as high as a factor of 2 (see Figure 2 in Duras et al., 2020).